OpenStudy (anonymous):

How many four-digit numbers can be formed using the digits 0, 1, 2, 3, 4, 5, 6, 7, 8, and 9 if the first digit cannot be 0? Repeated digits are allowed.

OpenStudy (anonymous):

So first digit we have 9 choice second digit 10 3rd 10 4th 10 9*10*10*10=9000

OpenStudy (anonymous):

why is the 1st digit 9?

OpenStudy (anonymous):

you have 9 choices 1 to 9

OpenStudy (anonymous):

ohh okay and why are the 2ns 3rd and 4th numbers 10?

OpenStudy (anonymous):

0- 9 ,----> 10 choices

OpenStudy (gw2011):

The first digit can be any one of 9 digits since 0 can't be the first digit. The other 3 digits can be any one of 10 digits--0 through 9. Therefore, you get: (9)(10^3)=(9)(1000)=9000

OpenStudy (anonymous):

gotcha!!!! thank you!!!

OpenStudy (gw2011):

You're welcome